# On GRPO Collapse in Search-R1: The Lazy Likelihood-Displacement Death Spiral

### v1

@@
Abstract
Tool-integrated (TI) reinforcement learning (RL) enables large language models (LLMs) to perform multi-step reasoning by interacting with external tools such as search engines and retrievers. Group Relative Policy Optimization (GRPO), exemplified by the recent Search-R1 (jin2025search), offers fast convergence and a value-free formulation that makes it appealing for this setting, yet consistently suffers from training collapse. We identify Lazy Likelihood Displacement (LLD), a systematic reduction or stagnation in the likelihood of both correct and incorrect responses, as the core mechanism driving this failure. LLD emerges early and triggers a self-reinforcing LLD Death Spiral, where declining likelihood leads to low-confidence responses, inflating gradients and ultimately causing collapse. We empirically characterize this process across models on a Search-R1-style, search-integrated question answering task, revealing a consistent three-phase trajectory: early stagnation, steady decay, and accelerated collapse. To address this, we propose a lightweight likelihood-preserving regularization $\rm LLDS$ for GRPO that activates only when a trajectory’s likelihood decreases, and regularize only the tokens responsible. This fine-grained structure mitigates LLD with minimal interference to optimization. Across seven open-domain and multi-hop QA benchmarks, our method stabilizes training, prevents gradient explosion, and yields substantial performance improvements, including +37.8% gains on Qwen2.5-3B and +32.0% gains on Qwen2.5-7B. Our results establish LLD as a fundamental bottleneck in GRPO-based TIRL and provide a practical path toward stable, scalable training of tool-integrated LLM.
@@ 1 Introduction
Figure 1: Comparative performance of LLDS and baseline methods on benchmark datasets. All baselines are built upon Qwen2.5-7B-Instruct. See Sec.˜6.1 for details.
@@
Large language models (LLMs) increasingly leverage external tools, such as search engines (jin2025search) and code-execution environments (feng2025retool; zeng2025simplerl), to augment their reasoning capabilities. This tool-integrated reasoning (TIR) paradigm has driven recent progress across factual question answering (jin2025search), image-based reasoning (jiang2025verltool), and mathematical problem solving (feng2025retool; jiang2025verltool). By enabling models to iteratively query tools, verify intermediate steps, and refine their thoughts, tool calls substantially elevate reasoning quality (qin2023toolllm). These advances naturally motivate the use of reinforcement learning (RL) to train LLMs to plan, interact with tools, and master multi-step decision making, as instantiated by recent GRPO-based frameworks such as Search-R1 (jin2025search).
@@
However, extending TIR to the RL setting introduces severe and persistent training instability. In particular, GRPO training in Search-R1-style jin2025search, search-integrated QA pipelines frequently exhibit sudden reward drops and catastrophic collapse (jin2025search; sun2025zerosearch). These failures are especially pronounced in multi-turn settings (xue2025simpletir), where tool feedback becomes part of the model’s conditioning context, creating long, highly entangled trajectories. Although prior work has observed these failures, the underlying mechanisms remain poorly understood.
@@
In this work, we firstly identify Lazy Likelihood Displacement (LLD) (deng2025effect; razin2024unintentional; ren2024learning), the stagnation or reduction of likelihood for both correct and incorrect responses during GRPO optimization (deng2025effect), as the fundamental but previously overlooked source of collapse in Search-R1–style tool-integrated reinforcement learning (TIRL). Using search-integrated question answering as a case study (see Fig.˜2), we show that LLD emerges early and persistently: even as rewards increase, the likelihood of correct responses enters a monotonic decline. This behavior appears across model scales, indicating that LLD is a structural failure mode of GRPO-based TIRL rather than a configuration-specific artifact. We further demonstrate that LLD drives a self-reinforcing LLD death spiral, where the resulting low-confidence regime amplifies negative-gradient influence from incorrect trajectories, accelerating likelihood decay, triggering entropy spikes, inflating likelihood ratios, and ultimately causing the large-gradient instability that leads to collapse.
@@
To counteract this failure mode, we propose a simple yet effective likelihood-preserving regularization $\rm LLDS$ that prevents harmful likelihood reductions. Our method integrates seamlessly with GRPO and introduces two layers of selectivity: (i) response-level gating, which activates the regularization only when a trajectory’s overall likelihood decreases, and (ii) token-level selectivity, which penalizes only the tokens responsible for the decrease. This fine-grained design directly mitigates LLD while minimally interfering with GRPO’s optimization behavior. By preventing unintentional downward likelihood drift, $\rm LLDS$ keeps optimization away from the unstable high-gradient regime and maintains healthy training dynamics. Empirically, this translates into stable training, suppressed gradient explosion, and consistent gains across seven open-domain and multi-hop QA benchmarks. Our main contributions are threefold:
  • We conduct the first systematic investigation of LLD in GRPO-driven TIRL and show that it arises with striking frequency, consistently progressing through a characteristic failure trajectory: steady likelihood decay, followed by gradient amplification and entropy explosion, and ultimately collapse.
@@
  • We propose a lightweight likelihood-preserving regularization $\rm LLDS$ that selectively regularize likelihood reductions and resolved the collapse issue in GRPO training.
@@
  • By stabilizing training, we realize significant performance gains across seven QA benchmarks (see Fig.˜1), demonstrating a more robust and reliable approach to RL-driven TIR.
Figure 2: We illustrate the likelihood displacement in tool-integrated RL training. The steady-decay phase (60-120) emerges when the reward begins to increase only gradually. In the subsequent acceleration phase (after step 120), the likelihood of correct responses drops sharply, accompanied by a sudden surge in gradient magnitude (red star), leading to gradient explosion. A zoomed-in view of the acceleration region further highlights this effect, showing a clearer likelihood displacement, where the gradient accelerates rapidly while the reward starts to decline.
@@ 2 Related Work
Tool-Integrated Reasoning and Agentic LLMs. Tool use has emerged as a powerful paradigm for equipping LLMs with adaptive reasoning capabilities. Early approaches relied on prompt-based orchestration (lu2023chameleon; shen2023hugginggpt) or multi-agent delegation frameworks to invoke tools without explicit training. Instruction-tuned models (gou2023tora; qin2023toolllm) later introduced structured tool-calling behaviors via supervised learning, but these systems remained largely static and were constrained to single-turn interactions. More recent work demonstrates that reinforcement learning can substantially enhance tool integration by enabling models to learn tool-usage policies through environmental feedback and task success. Notable systems such as RETool (feng2025retool), VERL-Tool (jiang2025verltool), and agentic LLM frameworks (maiagentic) support multi-step reasoning through dynamic tool use, self-verification, and error correction. This transition from static instruction-following to feedback-driven optimization has proven effective across a range of domains, including mathematical problem solving with code execution (xue2025simpletir), open-domain question answering with retrieval (jin2025search), SQL generation from natural language (jiang2025verltool), and multi-modal visual reasoning (gao2024multi).
Training Collapse in Tool-Integrated GRPO. GRPO (guo2025deepseek) has gained popularity in reinforcement learning due to its value-free, outcome-driven formulation. However, applying GRPO to train LLMs for multi-turn tool-integrated reasoning remains highly unstable (jin2025search). When initialized from base models using only verifiable rewards (Zero RL), GRPO-trained models often experience catastrophic failure, marked by abrupt reward collapse (xue2025simpletir; maiagentic; jin2025search). In contrast, Proximal Policy Optimization (PPO) (schulman2017proximal) tends to exhibit more stable behavior under comparable settings (jin2025search). These collapse dynamics have been consistently observed across various systems, including Search-R1 (jin2025search), SimpleTIR (xue2025simpletir), and ZeroSearch (sun2025zerosearch). While prior work has largely focused on empirical observations of failure, only xue2025simpletir attempt to explain the phenomenon, attributing it to low-likelihood incorrect responses, which inflate importance weights and destabilize learning. Yet this explanation does not fully account for the structural causes of such low-likelihood trajectories. Notably, we observe that reward degradation often precedes the final collapse, suggesting deeper optimization issues. In this work, we identify Lazy Likelihood Displacement (LLD) as the core mechanism underlying GRPO’s failure mode in multi-turn TIRL.
3 Priliminary
TIRL enables an LLM to interact with the external tools. We denote the query by $\bm{x}$, the tool feedback by $\bm{o}$, and the model’s action by $\bm{y}$. For a given query $\bm{x}$, the model autoregressively generates an action at turn $t$ according to
@@ 3 Priliminary
where $\bm{y}_{t}$ is produced based on the accumulated context, and $\bm{o}_{t-1}\sim\mathcal{T}$ denotes the tool feedback returned at turn t1t-1 if the previous action $\bm{y}_{t-1}$ invoked a tool $\mathcal{T}$ (e.g., issuing a search query). If no tool is invoked, $\bm{o}_{t-1}$ is an empty string. The multi-turn setting therefore corresponds to trajectories that involve multiple tool calls and, consequently, multiple rounds of tool feedback. However, tool feedback $\mathbf{o}_{t}$ is inherently out-of-distribution (OOD) for the pretrained language model, since such feedback originates from external environments rather than the model’s own generative distribution. Directly forcing the model to fit these feedback tokens introduces a large training–inference mismatch and may cause the model to memorize environment-specific feedback. Therefore, previous works commonly mask out feedback tokens during supervised training to stabilize optimization and avoid overfitting to external content (jin2025search).
Tool-Integrated GRPO with Feedback Mask.
The GRPO loss, introduced in DeepSeek-Math (shao2024deepseekmath) and DeepSeek-R1 (guo2025deepseek), enhances policy optimization by refining how reward and loss signals are computed. Unlike Proximal Policy Optimization (PPO) (schulman2017proximal), GRPO removes the need for value function estimation and instead uses group-relative reward normalization for optimization. Specifically, for a query–answer pair $(\bm{x},\bm{a})$, the policy $\pi_{\theta}$ samples $G$ responses:
@@
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}(\bm{x},\bm{a})\sim\mathcal{D}\\ \{\bm{y}_{i,0},\ldots,\bm{y}_{i,t},\bm{o}_{i,t},\bm{y}_{i,T_{i}}\}\sim(\pi_{\theta_{\mathrm{old}},\mathcal{T}})\end{subarray}}$
$\Bigg[\frac{1}{\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}|\hat{\bm{y}}_{i,t}|}\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}\sum_{k=1}^{|\hat{\bm{y}}_{i,t}|}\min\!\left(\gamma_{i,t,k}(\theta)\,\hat{A}_{i,t,k},\;\hat{A}_{i,t,k}\,\mathrm{clip}\big(\gamma_{i,t,k}(\theta),\,1-\varepsilon,\,1+\varepsilon\big)\right)\Bigg],$ (1)
where $\varepsilon$ is the clipping parameter, the likelihood ratio is defined as $\gamma_{i,k}(\theta)=\frac{\pi_{\theta}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}{\pi_{\theta_{\mathrm{old}}}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}$. Although feedback tokens $\bm{o}$ are masked out during loss computation, they still influence the context of subsequent token predictions.
@@ 4 Lazy Likelihood Displacement in Tool-integrated GRPO
Recent work (deng2025effect) introduces Lazy Likelihood Displacement (LLD) in GRPO for text-based, non-tool settings, showing that the likelihood of correct responses often decreases or improves only marginally during optimization. In this section, we extend LLD to the tool-integrated RL regime and define:
Definition 4.1 ( Tool-Lazy Likelihood Displacement)
Let $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{fin}}}$ denote the initial and finetuned policies obtained before and after optimizing a preference-learning objective $\mathcal{J}$ (e.g., Sec.˜3) over a dataset $\mathcal{D}$, with $\mathcal{J}(\theta_{\mathrm{fin}})<\mathcal{J}(\theta_{\mathrm{old}})$. Consider a tool-integrated trajectory consisting of alternating actions and feedback, $(\bm{y}_{0},\bm{o}_{0},\;\bm{y}_{1},\bm{o}_{1},\;\ldots,\;\bm{y}_{T}),$ where only the actions $\{\bm{y}_{t}\}_{t=0}^{T}$ are used in likelihood computation (feedback is masked). For each response action $\bm{y}_{t}$, define its log-likelihood change as
$\Delta_{t}(\mathbf{x},\bm{y}_{t}):=\ln\pi_{\theta_{\mathrm{fin}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t})-\ln\pi_{\theta_{\mathrm{old}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t}).$
We say that Lazy Likelihood Displacement (LLD) occurs for the action $t$ if $\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon$, where $\epsilon$ is a small or non-positive constant. And LLD occurs for the whole response if
@@
Thus, LLD captures the failure mode where one or more response actions exhibit negligible or even negative improvement in likelihood under the optimized policy.
To better understand the origin of LLD, we highlight how negative gradients suppress the likelihood of correct actions. The following informal theorem (the formal version and proof see App.˜A) restates Theorem 4.4 from deng2025effect in the setting of tool-integrated RL:
Theorem 4.2 (Informal: Trajectory-Level LLD in Tool-Integrated GRPO)
In tool-integrated GRPO, the likelihood of an otherwise correct response can decrease when (i) incorrect responses of low likelihood and (ii) incorrect responses whose embeddings closely resemble the correct one induce large negative gradients that dominate the positive updates. These forces jointly lead to trajectory-level Lazy Likelihood Displacement.
In the setting of tool-integrated GRPO, we observed with striking frequency that the likelihood of correct responses fails to improve. This indicates an especially acute form of LLD ($\epsilon\leq 0$), causing a progressive decay in the model’s overall output likelihood. This compounding decay is a distinctive failure mode of Search-R1 style tool-integrated RL.
5 Lazy Likelihood Displacement in tool-integrated GRPO
In this section, we illustrate the prevalence of LLD in tool-integrated GRPO training and show how the LLD ($\epsilon\leq 0$) can catastrophically drive the model into training collapse.

### v2

# On Group Relative Policy Optimization Collapse in Agent Search: The Lazy Likelihood-Displacement

@@
Abstract. Tool-integrated (TI) reinforcement learning (RL) enables large language models (LLMs) to perform multi-step reasoning by interacting with external tools such as search engines. Group Relative Policy Optimization (GRPO), exemplified by the recent Search-R1 (Jin et al., 2025), offers fast convergence and a value-free formulation that makes it appealing for this setting, yet consistently suffers from training collapse. We identify Lazy Likelihood Displacement (LLD), a systematic reduction or stagnation in the likelihood of correct response trajectories, as the core mechanism driving this failure. LLD emerges early and triggers a self-reinforcing LLD Death Spiral, where declining likelihood leads to low-confidence responses, inflating gradients and ultimately causing collapse. We empirically characterize this process across models on a search-integrated question answering task, revealing a three-phase trajectory: early stagnation, steady decay, and accelerated collapse. To address this, we propose a likelihood-preserving regularization $\rm LLDS$ that activates only when a response action’s likelihood decreases, and regularizes only the tokens responsible. This fine-grained structure mitigates LLD with minimal interference. Our method stabilizes training, prevents gradient explosion, and yields substantial performance improvements across seven benchmarks, including relative improvements of +45.2% on Qwen2.5-3B and +37.1% on Qwen2.5-7B over vanilla GRPO training. Our results establish LLD as a previously overlooked bottleneck in GRPO-based TIRL and provide a practical path toward stable, scalable training of tool-integrated RL.
Code     [Uncaptioned image] Models
@@ 1 Introduction
Figure 1: Comparative performance of LLDS and baseline methods on benchmark datasets. All baselines are built upon Qwen2.5-7B-Instruct. See Section˜6.1 for details.
@@
Large language models (LLMs) increasingly leverage external tools, such as search engines (Jin et al., 2025) and code-execution environments (Feng et al., 2025; Zeng et al., 2025), to augment their reasoning capabilities. This tool-integrated reasoning (TIR) paradigm has driven recent progress across factual question answering (QA) (Jin et al., 2025), image-based reasoning (Jiang et al., 2025), and mathematical problem solving (Feng et al., 2025; Jiang et al., 2025). By enabling models to iteratively query tools, tool calls substantially elevate reasoning quality (Qin et al., 2023). These advances naturally motivate the use of reinforcement learning (RL) to train LLMs to plan, interact with tools, and master multi-step decision making, as instantiated by TIRL frameworks such as Search-R1 (Jin et al., 2025)
@@
However, extending TIR to the RL setting leads to severe and persistent training instabilities. In particular, GRPO training in Search-R1–style pipelines frequently suffers from abrupt reward drops and catastrophic collapse (Jin et al., 2025; Sun et al., 2025). These failures are especially pronounced in multi-turn tool scenarios (Xue et al., 2025), where out-of-distribution tool feedback is incorporated into the model’s context. Although prior work observed these failures, the underlying mechanisms lack understanding.
@@
In this work, we first identify Lazy Likelihood Displacement (LLD) (Deng et al., 2025; Razin et al., 2024; Ren and Sutherland, 2024), the stagnation or reduction of likelihood for both correct and incorrect responses during GRPO optimization (Deng et al., 2025), as the fundamental but previously overlooked source of collapse in TIRL. Using search-integrated QA as a case study (see Figure˜2), we show that LLD emerges early and persistently: even as rewards increase, the likelihood of correct responses enters a monotonic decline. This behavior appears across model scales, indicating that LLD is a structural failure mode of GRPO-based TIRL rather than a configuration-specific artifact. We further demonstrate that LLD drives a self-reinforcing LLD death spiral, where the resulting low-confidence regime amplifies negative-gradient influence from incorrect trajectories, accelerating likelihood decay, triggering entropy spikes, inflating likelihood ratios, and ultimately causing the large-gradient instability that leads to collapse.
@@
To counteract this failure mode, we propose a simple yet effective LLD suppression ($\rm LLDS$) regularization that prevents harmful likelihood reductions. Our method integrates seamlessly with GRPO and introduces two layers of selectivity: (i) action-level gating, which activates the regularization only when a response action’s overall likelihood decreases, and (ii) token-level selectivity, which penalizes only the tokens responsible for the decrease. This fine-grained design directly mitigates LLD while minimally interfering with GRPO’s optimization behavior. By preventing unintentional downward likelihood drift, $\rm LLDS$ maintains stable training, suppressed gradient explosion, and consistent gains across seven open-domain and multi-hop QA benchmarks. Our main contributions are threefold:
$\bullet$ We identify LLD as a previously unrecognized failure mode in GRPO-based tool-integrated reinforcement learning, and show that it arises pervasively and follows a consistent, self-reinforcing trajectory of sustained likelihood decay, gradient amplification, entropy explosion, and eventual training collapse.
@@
$\bullet$ We propose a lightweight regularization $\rm LLDS$ that selectively regularize likelihood reductions and resolves the collapse issue in GRPO training.
@@
$\bullet$ By stabilizing training, we realize significant performance gains across QA benchmarks (see Figure˜1), demonstrating a more robust and reliable approach to TIRL.
@@ 2 Related Work
Tool-Integrated Reasoning and Agentic LLMs. Tool use enables LLMs to perform adaptive, multi-step reasoning. Early methods relied on prompt-based (Lu et al., 2023; Shen et al., 2023) or supervised tool calling (Gou et al., 2023; Qin et al., 2023), while recent RL-based systems (e.g., RETool (Feng et al., 2025), VERL-Tool (Jiang et al., 2025)) learn tool-usage policies from environmental feedback, achieving strong performance across QA (Jin et al., 2025), code-based math reasoning (Xue et al., 2025), SQL generation (Jiang et al., 2025), and multimodal tasks (Gao et al., 2024).
Training Collapse in Tool-Integrated GRPO. GRPO (Guo et al., 2025) is a value-free, outcome-driven RL method but exhibits severe instability in multi-turn tool-integrated settings, often collapsing when trained from base models with only verifiable rewards (Jin et al., 2025). This behavior is consistently observed in systems such as Search-R1 (Jin et al., 2025), SimpleTIR (Xue et al., 2025), and ZeroSearch (Sun et al., 2025), whereas PPO remains comparatively stable. Prior explanations attribute the collapse to training–inference mismatch (Liu et al., 2025) or low-likelihood incorrect responses (Xue et al., 2025), but do not explain PPO’s robustness or the structural origin of the instability. We identify Lazy Likelihood Displacement (LLD) as the underlying mechanism driving collapse in multi-turn TIRL (Section B.1). Recent works mitigate this via turn-level reward shaping, either relying on ground-truth documents (Zheng et al., 2025b) or an external LLM judge (Zhang et al., 2025), both introducing additional supervision or system cost and increased vulnerability to reward hacking (Guo et al., 2025). Recently, TreeGRPO (Ji et al., 2025) provides dense reward using tree-based rollout, but increases rollout and computational overhead. In contrast, our approach shows that outcome rewards alone suffice to achieve superior performance once GRPO is stabilized. Detailed related work see Section˜B.1.
3 Preliminary
We denote the query by $\bm{x}$, the tool feedback by $\bm{o}$, and the model’s action by $\bm{y}$. For a given query $\bm{x}$, the model autoregressively generates an action at turn $t$ according to
@@ 3 Preliminary
where $\bm{y}_{t}$ is produced based on the accumulated context, and $\bm{o}_{t-1}\sim\mathcal{T}$ denotes the tool feedback returned at turn t1t-1 if the previous action $\bm{y}_{t-1}$ invoked a tool $\mathcal{T}$ (e.g., issuing a search query). If no tool is invoked, $\bm{o}_{t-1}$ is an empty string. The multi-turn setting thus corresponds to trajectories involving multiple tool calls. However, tool feedback $\mathbf{o}_{t}$ is inherently out-of-distribution (OOD) for the LLM, as it originates from external environments rather than the model’s generative distribution. Thus, prior work (Jin et al., 2025) masks feedback tokens during training to stabilize optimization.
Tool-Integrated GRPO with Feedback Mask. GRPO, introduced in DeepSeek-Math (Shao et al., 2024) and DeepSeek-R1 (Guo et al., 2025) and later applied to tool use (Jin et al., 2025), is a value-free policy optimization method with group-relative reward normalization. Specifically, for a query–answer pair $(\bm{x},\bm{a})$, the policy $\pi_{\theta}$ samples $G$ responses:
@@
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}(\bm{x},\bm{a})\sim\mathcal{D}\\ \{(\bm{y}_{i,t},\bm{o}_{i,t})_{t=0}^{T_{i}-1},\bm{y}_{i,T_{i}}\}\sim(\pi_{\theta_{\mathrm{old}},\mathcal{T}})\end{subarray}}$
$\Bigg[\frac{1}{\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}|\hat{\bm{y}}_{i,t}|}\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}\sum_{k=1}^{|\hat{\bm{y}}_{i,t}|}\min\!\Bigl(\gamma_{i,t,k}(\theta)\,\hat{A}_{i,t,k},\;$
$\hat{A}_{i,t,k}\,\mathrm{clip}\big(\gamma_{i,t,k}(\theta),\,1-\varepsilon,\,1+\varepsilon\big)\Bigr)\Bigg],$ (1)
where $\varepsilon$ is the clipping parameter, the likelihood ratio is defined as $\gamma_{i,t,k}(\theta)=\frac{\pi_{\theta}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}{\pi_{\theta_{\mathrm{old}}}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}$. Although feedback tokens $\bm{o}$ are masked out, they still influence the context of subsequent token predictions.
Figure 2: We illustrate the likelihood displacement in tool-integrated RL training. The steady-decay phase (60-120) emerges even when the reward increases. In the subsequent acceleration phase (after step 120), the likelihood of correct responses drops sharply, accompanied by a sudden surge in gradient magnitude (red star), leading to gradient explosion. A zoomed-in view of the acceleration region further highlights this effect, showing a clearer likelihood displacement, where the gradient accelerates rapidly while the reward starts to decline.
@@ 4 Lazy Likelihood Displacement in Tool-integrated GRPO
We identify a critical but previously unrecognized failure mode in tool-integrated GRPO, in which likelihood displacement progressively accelerates through compounding low-confidence responses and ultimately destabilizes training. While (Deng et al., 2025) identifies LLD as a training pathology in single-turn GRPO, their analysis is restricted to closed-loop text generation without external observations. In contrast, tool-integrated reinforcement learning fundamentally changes the learning dynamics: LLD becomes a trajectory-level instability that interacts with out-of-distribution tool feedback and prefix-shared decision branches, leading to systematic likelihood decay, gradient amplification, and eventual training collapse.
We further show that this phenomenon arises from two structural factors unique to the tool-integrated GRPO setting (analyzed in Section 5.4) and consistently leads to catastrophic collapse behavior. We formalize and define LLD in the tool-integrated RL regime in this section.
Definition 4.1 (Tool-LLD).
Let $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{fin}}}$ denote the initial and finetuned policies obtained before and after optimizing the objective $\mathcal{J}$ (e.g., Section˜3) over a dataset $\mathcal{D}$, with $\mathcal{J}(\theta_{\mathrm{fin}})<\mathcal{J}(\theta_{\mathrm{old}})$. Consider a tool-integrated trajectory consisting of alternating actions and feedback, $(\bm{y}_{0},\bm{o}_{0},\;\bm{y}_{1},\bm{o}_{1},\;\ldots,\;\bm{y}_{T}),$ where only the actions $\{\bm{y}_{t}\}_{t=0}^{T}$ are used in likelihood computation (feedback is masked). For each response action $\bm{y}_{t}$, define its log-likelihood change as
$\Delta_{t}(\mathbf{x},\bm{y}_{t})$ $:=\ln\pi_{\theta_{\mathrm{fin}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t})$
$-\ln\pi_{\theta_{\mathrm{old}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t}).$
We say that LLD occurs for the action $t$ if $\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon$, where $\epsilon$ is a small or non-positive constant. And LLD occurs for the whole response if
@@
Therefore, LLD accounts for the failure scenario where even correct response actions fail to boost confidence (or even reduce it) when such an enhancement is explicitly required by the loss function.
(a) Likelihood displacement across training steps.
(b) Mechanism
Figure 3: (a) Effect of likelihood displacement across different training iterations for the Qwen2.5-3B-Instruct model. Results are computed on the first 50 samples of the training set, discarding cases where all responses are uniformly correct or uniformly incorrect. Bars below zero (orange) indicate samples whose correct responses’ likelihood decreases after training. (b) Illustration of the mechanism: as training progresses, the likelihood of correct responses decreases while hidden-state similarity increases, leading to stronger LLDS.
We generalize the negative-gradient characterization of group-relative updates in (Deng et al., 2025) to a multi-turn setting with external observations, where the learning dynamics differ qualitatively. We present the following informal theorem (Formal statement and proof in Appendix˜A).
Theorem 4.2 (Informal: Trajectory-Level LLD in Tool-Integrated GRPO).
In tool-integrated GRPO, the likelihood of a correct response action can decrease when incorrect responses satisfy two conditions: (i) Low likelihood: The model assigns low probability to incorrect responses, causing large prediction-error weights that magnify their influence, and (ii) High embedding similarity: Incorrect responses share similar representations with correct ones, causing negative gradients to interfere with positive updates. As they frequently do in tool-integrated settings (see Section 5.4), negative gradients dominate, causing LLD.
In the setting of tool-integrated GRPO, we observed with striking frequency that the likelihood of correct responses reduces (Figure˜2). This indicates an especially acute form of LLD ($\epsilon\leq 0$), causing a progressive decay in the model’s output likelihood. This compounding decay is a distinctive failure mode of Search-R1 style TIRL.
5 LLD in Tool-Integrated GRPO Collapse
This section examines the prevalence of LLD in tool-integrated GRPO and illustrates how sustained likelihood decay ($\epsilon\leq 0$) leads to catastrophic training collapse. All analyses in this section use the NQ dataset (Kwiatkowski et al., 2019) unless otherwise stated.